A pole support by a steel wire is in the way of construction site. In order to move heavy to machine, the steel wire must be moved 4m closer to the pole. If the pole height is 30m and the cable is currently fastened to the ground 12m from the pole, how much steel wire need to cut off from the cable when they move it?
1
Expert's answer
2019-10-10T09:51:14-0400

For the given problem the length of steel wire will be the hypotenuse of triangle with base 12 metre and height 30 metre.

So,

Initial length of wire

"=\sqrt{30^2+12^2}=32.311\ m"

Finally base will be

"12-4\\=8\ m"

So,

Final length of wire

"=\sqrt{30^2+8^2}=31.048\ m"

Length which need to cut = initial length - final length

"=32.311-31.048=1.263\ m"